OpenStudy (anonymous):

express this polynomial in standard form please 4x(2x^2 - 2x + 5) -6x^2(2x^2-7x -2)

OpenStudy (amistre64):

just multiply them out and add them up ..

OpenStudy (anonymous):

if it was that easy I would not ask, please assist if you can

OpenStudy (anonymous):

\[-12 x^4+50 x^3+4 x^2+20 x\]

OpenStudy (anonymous):

but that is not an assist, that is just an answer. lets do it step by step

OpenStudy (amistre64):

oh but it is that easy .... what part is giving you troubles?

OpenStudy (anonymous):

\[4x(2x^2-2x+5)=4x\times 2x^2+4x\times -2x+4x\times 5\]

OpenStudy (anonymous):

that is the first term via the distributive law. multiply everything by 4x

OpenStudy (amistre64):

seeing how they dont hang around to work thru it ... its hard to "help"

OpenStudy (anonymous):

now we multuiply and add exponents to get \[8x^3-8x^2+20x\]

OpenStudy (anonymous):

really? then off i go

OpenStudy (anonymous):

ok that was great, be patient man I am an old man trying to catch up with some algebra. I got it though thanks

OpenStudy (anonymous):

yw